OpenStudy (anonymous):

can some one explain to me how to simplify 18y^3/36xy^2

OpenStudy (kirbykirby):

18/36 = 1/2 y^3/(xy^2) = y/x... just subtract the exponents of the y's (3-2 = 1), and just keep the x in the denominator hence, 1/2* (y/x) = y/(2x)